OpenStudy (anonymous):

plz help to solve this integration: d(2x-1)/x * sqrt(x^2-2)

OpenStudy (anonymous):

\[\frac{ d(2x-1) }{ x } \times \sqrt{x^2-2}\] ?

OpenStudy (anonymous):

no.. the sqrt portion is multiplied with x i.e. its bellow the bar.. x * sqrt(x^2-2)

OpenStudy (anonymous):

plz sum1 solve it.. urgent!!!\[d(2x-1)\x \sqrt{(x ^{2}-2)}{ ? }{ ? }\]